### v1

# Learning Laplacian Eigenvectors: a Pre-training Method for Graph Neural Networks

Abstract
We propose a novel framework for pre-training Graph Neural Networks (GNNs) by inductively learning Laplacian eigenvectors. Traditional Message Passing Neural Networks (MPNNs) often struggle to capture global and regional graph structure due to over-smoothing risk as network depth increases. Because the low-frequency eigenvectors of the graph Laplacian matrix encode global information, pre-training GNNs to predict these eigenvectors encourages the network to naturally learn large-scale structural patterns over each graph. Empirically, we show that models pre-trained via our framework outperform baseline models on a variety of graph structure-based tasks. While most existing pre-training methods focus on domain-specific tasks like node or edge feature reconstruction, our self-supervised pre-training framework is structure-based and highly flexible. Eigenvector-learning can be applied to all graph-based datasets, and can be used with synthetic features when task-specific data is sparse.
1 Introduction
Graph Neural Networks (GNNs) have become a powerful tool in node and graph representation learning, with successful applications across domains ranging from biomedicine [1, 2, 3, 4] to social networks [5]. However, traditional message passing GNNs often struggle to capture the global and regional structure of the graph, primarily due to the risk of over-smoothing as network depth increases [6, 7]. One promising solution is to pre-train GNNs on structure-aware tasks [4]. In particular, pre-training a GNN to produce the low-frequency eigenvectors of the graph Laplacian enables it to learn global patterns within the data. These low-frequency eigenvectors capture global and regional structure; by training the GNN to produce the lower frequency components, we encourage it to focus on the larger-scale structure of the graph. The pre-trained model can then be fine-tuned on other downstream, context-aware tasks, or its embeddings can be used as positional encodings for Graph Transformer networks. Moreover, since Laplacian eigenvector learning is entirely structure-based, it can be used as a training task for transfer learning and graph foundation models.
While some structure-informed pre-training tasks have been proposed, the majority are tailored to specific domains and rely on access to informative node features [3, 2]. Common strategies include context-prediction and motif or cluster prediction [3, 4, 2]. Existing domain-agnostic pre-training strategies also suffer from over-smoothing, as they rely solely on deep message passing GNNs [3, 1, 2] with basic node-wise prediction heads. This highlights the need for an alternative prediction head that can naturally capture and synthesize long-range structural information.
Our main contributions are as follows:
  1. 1.
    We introduce Laplacian eigenvector-learning as a structure-based pre-training task for GNNs.
  2. 2.
    We introduce a global MLP head that enables long-range interaction between vertices within the graph, improving the effectiveness of eigenvector-based pre-training.
  3. 3.
    We demonstrate that our pre-training framework provides performance improvements over baseline models on various structure-based downstream applications.
2 Background and Notation
The Laplacian $L$ of a graph $G$ is defined as:
L=DAL=D-A
where $D$ is the diagonal degree matrix and $A$ is the unnormalized adjacency matrix of $G$.
Let $\lambda_{1},\lambda_{2},\dots\lambda_{k}$ denote the $k$ lowest eigenvalues of $L$ in nondecreasing order. Let $\psi_{1},\psi_{2},\dots\psi_{k}$ denote the corresponding eigenvectors.
By Courant-Fischer, the eigenvectors of $L$ (and the eigenvectors of any Hermitian matrix) can be expressed as the following iterative optimization problem:
$\psi_{k}\in\operatorname*{argmin}_{\begin{subarray}{c}\|x\|=1\\ x\perp\psi_{1},\dots,\psi_{k-1}\end{subarray}}x^{\top}Lx.$
The term $\frac{x^{\top}Lx}{x^{\top}x}$ is known as the Rayleigh quotient; because we normalize our predicted eigenvectors, we simply treat this as $x^{\top}Lx$.
3 Method
3.1 Model Architecture
Base GNN: The base GNN model takes in a graph with augmented node features and generates learned node representations via neighborhood message passing and update steps. Any GNN architecture may be selected as the base model to fit the needs of the dataset and downstream application.
Graph-level MLP: We concatenate the node-wise output of the base GNN model to form a graph-level aggregated representation. We then pass the aggregated vector through an MLP model to produce the low-frequency Laplacian eigenvectors. Concatenating the node embeddings prior to applying the MLP allows the model to learn relationships between distant nodes without risking over-smoothing.
Previous eigenvector-learning methods use a node-wise MLP head, processing each node’s eigencoordinates independently based on their learned hidden embedding [8, 9, 1].
3.2 Node Feature Augmentation
To provide the model with additional structural information, we augment node features with structure-based embeddings. We propose two kinds of embeddings: (1) wavelet positional embeddings, which encode relative positional information between nodes, and (2) diffused dirac embeddings, which encode local connectivity structures around each node. Both embeddings use the random walk matrix, and capture local aggregate information on each node. We discuss these embeddings in detail in A.4.
3.3 Loss function
We minimize a weighted sum of two loss functions: (1) eigenvector loss and (2) energy loss.
Let $\hat{U}$ denote a matrix of column vectors $\hat{u}_{i}$, where each $\hat{u}_{i}$ denotes the $i$th predicted eigenvector.
Eigenvector loss, used in [10], measures the difference between each $L\hat{u}_{i}$ and $\lambda_{i}\hat{u}_{i}$:
$\mathcal{L}_{eigvec}=\frac{1}{k}\lVert(L\hat{U}-\hat{U}\Lambda_{k})\rVert$
Energy loss, used in [8, 9, 11], aims to minimize the sum of Rayleigh quotients:
[…]
To ensure the model does not output $k$ copies of the trivial eigenvector, we impose orthogonality on the final outputs of the model via QR decomposition, as proposed in [8].
We show that our loss functions satisfy necessary sign and basis invariances. Full proofs can be found in A.3.
3.4 Pre-Training Algorithm
Algorithm 1 Eigenvector Learning
1:Graph $G=(V,E)$; augmented node features $\tilde{X}=\{\tilde{x}_{j}\}$; Base GNN
2:Output Pre-trained GNN model, $k$ lowest-frequency eigenvectors
3:for $i<\text{Pre-Training Epochs}$ do
[…]
5:  $\vec{Z}\leftarrow[\vec{z}_{1},\dots,\vec{z}_{n}]\in\mathbb{R}^{nd}$
6:  $\tilde{U}\leftarrow\textsc{MLP}(\vec{Z})$
[…]
9:  Back-propagate Loss, update model weights
10:end for
11:return BaseGNN
Figure 1: Visual illustration of Laplacian Eigenvector GNN pre-training framework.
4 Experimental Results
4.1 Standalone pre-training
We pre-train a standard Graph Isomorphism Network (GIN) [12] on the eigenvector-learning task with the graph-level MLP head. Once the model has been pre-trained, we replace the graph-level MLP head with a downstream prediction MLP and fine-tune model weights. We evaluate our pre-training framework on two molecular datasets, ZINC-12k [13] and QM9 [14]. We record results of our experiments in Table 1. On 7 of 8 total targets, eigenvector pre-training substantially improves downstream performance.
Table 1: Test MAE ($\downarrow$) performance comparison on ZINC (single metric) and QM9 (first seven target properties).
ZINC QM9 Model Penalized $\log p$ $\mu$ $\alpha$ $\varepsilon_{\text{HOMO}}$ $\varepsilon_{\text{LUMO}}$ $\Delta_{\varepsilon}$ $\braket{R^{2}}$ ZPVE Pre-Trained GIN 0.353 0.484 0.489 0.00353 0.00371 0.00513 28.103 0.000477 Baseline GIN 0.438 0.472 1.132 0.00386 0.00399 0.00562 50.909 0.002400
4.2 Augmenting existing pre-training method
We augment the existing molecular pre-training methods proposed in [3] with eigenvector-learning. In particular, [3] proposes node-level pre-training tasks (context prediction and masking) on ZINC15 [13], followed by a graph-level supervised pre-training task on ChEMBL [15, 16]. We augment the graph-level supervised pre-training step by adding an additional MLP head to predict eigenvectors, and we evaluate on five downstream datasets based on [4].
Detailed results are shown in Table 2. Eigenvector-learning consistently improves performance for the masking pre-training pipeline, but achieves mixed results on the context prediction pipeline. Notably, performance for the masking pipeline was increased for all five datasets when performing eigenvector pre-training with the graph-level MLP head.
Table 2: Test ROC-AUC (%, $\uparrow$) performance on 5 molecular prediction tasks when augmenting an existing pre-training method on a GIN base model. Sup. refers to the original supervised pre-training as implemented in [3], and Sup.+ refers to supervised training with additional eigenvector-learning (our method). Results for no pre-training are taken directly from [4]. All methods are tuned over seven learning rates and averaged over three seeds.
Dataset BACE BBBP Tox21 ToxCast SIDER Pretrain method MLP Head ContextPred, Sup.+ Graph-level $79.62\pm 3.63$ $\mathbf{70.76\pm 1.64}$ $77.94\pm 0.11$ $\mathbf{66.13\pm 0.34}$ $60.05\pm 0.99$ ContextPred, Sup.+ Node-wise $75.87\pm 3.11$ $68.74\pm 1.07$ $78.86\pm 0.06$ $63.78\pm 0.32$ $59.83\pm 0.53$ ContextPred, Sup. - $\mathbf{84.98\pm 1.28}$ $68.25\pm 0.48$ $77.44\pm 0.19$ $64.01\pm 0.81$ $\mathbf{62.87\pm 0.89}$ Masking, Sup.+ Graph-level $80.71\pm 3.84$ $68.33\pm 0.89$ $\mathit{79.09\pm 0.25}$ $\mathit{65.96\pm 0.20}$ $\mathit{62.41\pm 1.77}$ Masking, Sup.+ Node-wise $\mathit{81.02\pm 1.67}$ $\mathit{69.94\pm 1.76}$ $\mathbf{79.33\pm 0.41}$ $65.14\pm 0.44$ $59.38\pm 1.11$ Masking, Sup. - $75.42\pm 2.64$ $67.36\pm 4.60$ $78.33\pm 0.24$ $64.88\pm 0.82$ $61.6\pm 1.78$ No pre-training - $75.77\pm 4.29$ $69.62\pm 1.05$ $75.52\pm 0.67$ $63.67\pm 0.32$ $59.07\pm 1.13$
Table 3: Test MAE ($\downarrow$) performance on ZINC (12k subset) dataset when augmenting a graph structural encoder. All results using no structural encoder or the base GPSE are taken directly from [1]. GPSE+ refers to GPSE with modified eigenvector-learning (our method). Our experimental results are averaged over three seeds.
Base model Structural encoder MLP Head MAE GPS GPSE+ Graph-level $\mathbf{0.0629\pm 0.0016}$ GPSE+ Node-wise $0.0663\pm 0.0024$ GPSE - $\mathit{0.0648\pm 0.0030}$ None - $0.118\pm 0.005$ GIN GPSE+ Graph-level $0.1231\pm 0.0026$ GPSE+ Node-wise $0.1299\pm 0.0010$ GPSE - $0.124\pm 0.002$ None - $0.285\pm 0.004$
4.3 Augmenting existing graph structural encoder
We modify Graph Positional and Structural Encoder (GPSE) [1]. While GPSE already incorporates eigenvector-learning by including Laplacian positional encodings (LapPE) as a prediction target, GPSE uses MAE and cosine similarity loss on the absolute value of each eigenvector and a node-wise prediction head for every node property. We replace GPSE’s eigenvector-learning component with our own, using a separate MLP head and eigenvector loss. We keep all other GPSE model settings the same.
Following [1], we pre-train our modified GPSE model on MolPCBA [17]. We evaluate the effectiveness of these encodings by training both the Transformer model GPS [18] and a standard GIN, augmented with these encodings, on the downstream molecule property prediction task for the ZINC 12k [13] subset.
We report results on the effectiveness of these new learned encodings in Table 3. Modified eigenvector-learning with a graph-level MLP improves performance on downstream performance for both the GIN and GPS models, with our graph-level GPS+GPSE+ configuration achieving SOTA performance over all model and encoding configurations tested in [1]. We also demonstrate that our choice of loss function is crucial for eigenvector-learning in A.7.
5 Limitations and Future Work
There are several promising future directions toward improving the Laplacian eigenvector pre-training framework. We have demonstrated the effectiveness of the framework for pre-training and fine-tuning a GNN on the same dataset, or on domain-related datasets. However, we are yet to explore the effectiveness of eigenvector pre-training as a transfer learning framework.
Further, the practical implementation of the graph-level MLP requires adding padding to the concatenated node embeddings to accommodate for graphs of differing sizes. This creates an additional challenge for the MLP to learn meaningful relationships between the individual node embeddings within the graph-level vector. One could explore other ways of creating a rich graph-level representation while avoiding this challenge.
Appendix A Appendix
A.1 Full eigenvector pre-training pipeline
We provide a broad algorithmic outline of eigenvector pre-training process in Algorithm 2.
Algorithm 2 Structure-Informed Graph Pre-training Framework
1:Graph $G=(V,E)$; node features $X=\{x_{j}\}$; training labels $Y$; untrained Base GNN; untrained Downstream Prediction Head
2:Trained Base GNN and Downstream Prediction Head
[…]
4:BaseGNN $\leftarrow\textsc{EigvecPreTrain}(G,\tilde{X},\textsc{BaseGNN})$
5:for $i<\text{Fine-tuning Epochs}$ do
[…]
7:  $\vec{Z}\leftarrow[\vec{z}_{1},\dots,\vec{z}_{n}]$
8:  $\hat{Y}\leftarrow\textsc{DownstreamHead}(\vec{Z})$
9:  $\textit{Loss}=\textsc{LossCriterion}(\hat{Y},Y)$
10:  Backpropagate Loss, update model weights
11:end for
12:return BaseGNN, DownstreamHead
A.2 Loss function
Eigenvector loss, per-vector form:
$\mathcal{L}_{eigvec}=\frac{1}{k}\sum_{i=1}^{k}\lVert L\hat{u}_{i}-\lambda_{i}\hat{u}_{i}\rVert$
Eigenvector loss, matrix form:
$\mathcal{L}_{eigvec}=\frac{1}{k}\lVert(L\hat{U}-\hat{U}\Lambda_{k})\rVert$
Energy loss, per-vector form:
$\mathcal{L}_{energy}=\frac{1}{k}\sum_{i=1}^{k}\hat{u}_{i}^{\top}L\hat{u}_{i}$
Energy loss, matrix form:
[…]
Energy loss is order-invariant and rotation invariant (see A.3); for applications in clustering, this is reasonable. However, we would like the model to learn the eigenvectors in their specific order, so we also define absolute energy loss, matching the Rayleigh quotient with the ground-truth eigenvalue:
$\mathcal{L}_{energy\_abs}=\frac{1}{k}\sum_{i=1}^{k}\lvert\hat{u}_{i}^{\top}L\hat{u}_{i}-\lambda_{i}\rvert$
This can be written as, in matrix form:
$\mathcal{L}_{energy\_abs}=\frac{1}{k}\operatorname{Tr}\lvert\hat{U}^{\top}L\hat{U}-\Lambda_{k}\rvert$
In practice, we do not show any results using absolute energy loss, and instead linearly combine energy loss with eigenvector loss to avoid order and rotation invariance. However, absolute energy loss remains an interesting avenue to explore.
A.2.1 Orthogonality
To ensure the model does not output $k$ copies of the trivial eigenvector, we must give the model orthogonality constraints on the output vectors. There are again two reasonable choices here: (1) forced orthogonality and (2) orthogonality loss.
Forced orthogonality, used in [8], imposes orthogonality on the final outputs of the model via QR decomposition. In other words, if $\hat{U}^{\prime}$ is the initial output to the model, $Q$ is an $n\times k$ matrix with orthonormal columns, and $R$ is a $k\times k$ upper triangular matrix, then we achieve the final output $\hat{U}$ as such:
$QR=\hat{U}^{\prime}$
$\hat{U}=Q$
Orthogonality loss, used in [9, 11, 10] imposes a softer constraint, encouraging orthogonality by penalizing the model for producing pairwise similar vectors. This can be written as:
$\mathcal{L}_{ortho}=\frac{1}{k}\lVert\hat{U}^{\top}\hat{U}-I\rVert$
Based on preliminary testing, we found that forced orthogonality improved performance on the eigenvector-learning, and thus use forced orthogonality in all of our experiments.
A.3 Energy and eigenvector losses are sign and basis invariant
A.3.1 Definition of basis invariance
Consider any eigenspace spanned by ground truth eigenvectors $\psi_{j},\psi_{j+1},\dots\psi_{j+k-1}]=V$. Also recall that, by Spectral Theorem, we can decompose any vector $u$ into a linear combination of all eigenvectors:
$u=\sum_{i=1}^{n}c_{i}\psi_{i}$
Then a loss function is basis invariant if any rotation of the projected component $VV^{\top}u$ does not change the loss incurred by $u$. In other words, $u$ gets to arbitrarily “choose" with what basis it wants to express its $VV^{\top}u$ component. Sign invariance is a special case of basis invariance, where changing sign is equivalent to rotating over a one-dimensional subspace (note that this is slightly stronger than the most apparent form of sign invariance, where we would say $\mathcal{L}(u)=\mathcal{L}(-u)$; instead, we can flip any component $c_{i}\psi_{i}$ of $u$ when decomposed in terms of eigenvectors).
Definition 1 (Basis invariance).
Consider an eigenspace spanned by ground truth eigenvectors $\psi_{j},\psi_{j+1},\dots\psi_{j+k-1}]=\Psi\in\mathbb{R}^{n\times k}$. Consider an eigenspace rotation $R_{\Psi}$ defined as such:
$R_{\Psi}=\Psi A\Psi^{\top}+(I_{n}-\Psi\Psi^{\top}),A\in\textsc{SO}(k)$
A loss function $\mathcal{L}(u)$ is basis invariant if, for all such $\Psi,R_{\Psi}$, $u\in\mathbb{R}^{n}$, we have:
$L(u)=L(R_{\Psi}u)$
A.3.2 Proofs
Lemma 1 (Energy loss is basis invariant).
For any $R_{\Psi}$ and a single eigenvector prediction $u\in\mathbb{R}^{n}$, we have:
$u^{\top}Lu=(R_{\Psi}u)^{\top}L(R_{\Psi}u)$
Proof.
First note that $R_{\Psi}$ is orthogonal; the set of all $R_{\Psi}$ describes a subset of $SO(k)$ where only the $k$ basis vectors in $\Psi$ are rotated. Thus, we have $R_{\Psi}^{\top}R_{\Psi}=I$.
In addition, because $\Psi$ is an eigenspace, all columns are eigenvectors with a shared eigenvalue $\lambda$. Then we have:
$R_{\Psi}L=\Psi A\Psi^{\top}L+L-\Psi\Psi^{\top}L=\lambda\Psi A\Psi^{\top}+L-\lambda\Psi\Psi^{\top}=L\Psi A\Psi^{\top}+L-L\Psi\Psi^{\top}=LR_{\Psi}$
Then we have:
$R_{\Psi}^{\top}LR_{\Psi}=R_{\Psi}^{\top}R_{\Psi}L=L$
Thus, for any $u$, we have:
$u^{\top}Lu=u^{\top}R_{\Psi}^{\top}LR_{\Psi}u=(R_{\Psi}u)^{\top}L(R_{\Psi}u)$
Lemma 2 (Eigenvector loss is basis invariant).
For any $R_{\Psi}$ and a single eigenvector prediction $u\in\mathbb{R}^{n}$ and ground truth eigenvalue $\lambda$, we have:
$\lVert Lu-\lambda u\rVert=\lVert L(R_{\Psi}u)-\lambda(R_{\Psi}u)\rVert$
Proof.
We know, from our proof above in Lemma 1, that $R_{\Psi}L=LR_{\Psi}$. Because $R_{\Psi}\in\textsc{SO}(k)$, we have $\lVert R_{\Psi}x\rVert=\lVert x\rVert$ for any $x\in\mathbb{R}^{n}$. Then we have:
$\lVert Lu-\lambda u\rVert=\lVert R_{\Psi}(Lu-\lambda u)\rVert$
$\lVert Lu-\lambda u\rVert=\lVert L(R_{\Psi}u)-\lambda(R_{\Psi}u)\rVert$
We have an even stronger statement of invariance for energy loss: given a predicted set of $k$ orthogonal vectors, rotating the vectors within the same subspace does not impact loss. In other words, a model trained on energy loss only needs to predict the correct subspace of $k$ eigenvectors. This is clearly not true of eigenvector loss. Depending on the application, this kind of invariance can be good or bad.
Lemma 3 (Energy loss is rotation invariant).
Let $L$ be a Laplacian matrix and $V\subseteq\mathbb{R}^{n}$ be some $k$-dimensional subspace. Suppose $U=[u_{1},u_{2},\dots,u_{k}],W=[w_{1},w_{2},\dots,w_{k}]\in\mathbb{R}^{n\times k}$ are both orthonormal bases for $V$. Then we have:
$\frac{1}{k}\operatorname{Tr}(U^{\top}LU)=\frac{1}{k}\operatorname{Tr}(W^{\top}LW)$
Proof.
Note that $UU^{\top}=WW^{\top}$, as they are both orthogonal projectors for the same subspace. Then we have, by the cyclic property of trace:
$\frac{1}{k}\operatorname{Tr}(U^{\top}LU)=\frac{1}{k}\operatorname{Tr}(UU^{\top}L)=\frac{1}{k}\operatorname{Tr}(WW^{\top}L)=\frac{1}{k}\operatorname{Tr}(W^{\top}LW)$
A.4 Node feature augmentation
The diffusion operator $P$ of a graph $G$ is defined as:
$P=D^{-1}A$
Each entry $P_{ij}$ represents the probability of starting a random walk at node $i$ and ending at node $j$ after one step. One can also take powers of the diffusion operator, $P^{t}$. Each entry of the powered matrix, $P^{t}_{ij}$, represents the probability of starting a random walk at node $i$ and ending at node $j$ after t steps
The $j^{\text{th}}$ wavelet operator $\Psi_{j}$ of a graph $G$ is defined as:
$\Psi_{j}=P^{2^{j-1}}-P^{2^{j}}$
$\Psi_{0}=I-P$
A wavelet bank, $\mathcal{W}_{J}$ is a collection of wavelet operators such that:
$\mathcal{W}_{J}=\{\Psi_{j}\}_{0\leq j\leq J}\cup P^{2^{J}}$
Wavelet positional embeddings encode information about the relative position of each node within the graph. We randomly select two nodes from each graph, $i$ and $j$, and start dirac signals $\delta_{i},\delta_{j}$. We then apply these signals to each wavelet, $\Psi_{k}$, in our wavelet bank. The wavelet positional embedding for node $m$ is the $m^{\text{th}}$ row of the resulting matrix.
$w_{m,k}=\Psi_{k}(m,\cdot)\begin{bmatrix}|&|\\ \delta_{i}&\delta_{j}\\ |&|\end{bmatrix}$
$w_{m}=\begin{bmatrix}w_{m,1}&\dots&w_{m,J}\end{bmatrix}$
Diffused dirac embeddings encode information about the connectedness of each node and its neighbors. For each node, $m$, we apply the $m^{\text{th}}$ row of the diffusion matrix $P$ to each wavelet $\Psi_{k}$ in our wavelet bank. As above, the difused dirac embedding for node $m$ is the $m^{\text{th}}$ row of the resulting matrix.
$d_{m,k}=\Psi_{k}(m,\cdot)~P(m,\cdot)^{\top}$
$d_{m}=\begin{bmatrix}d_{m,1}&\dots&d_{m,J}\end{bmatrix}$
Lemma 4 (Uniqueness up to co-spectrality).
Let $G_{1},G_{2}$ be graphs of size $n$ with Laplacian matrices $L_{1},L_{2}$ respectively. Let $d^{1}_{m},d^{2}_{m}$ represent the diffused dirac embeddings for each node in $G_{1},G_{2}$. Then if $L_{1}$ and $L_{2}$ have different eigenvalues, $\{d^{1}_{m}:m\leq n\}\neq\{d_{m}^{2}:m\leq n\}$
Proof.
Consider the random-walk Laplacian of a graph: $L_{rw}:=I-D^{-1}A=I-P$. Moreover, note that $L_{rw}=D^{-1}L$. Observe that
$L_{rw}Dv$ $=D^{-1}LDv$
$=D^{-1}U\Lambda U^{\top}Dv$
[…]
Where $U=\begin{bmatrix}\psi_{1}&\dots&\psi_{n}\end{bmatrix}$, with $\psi_{i}$ orthonormal eigenvectors of $L$ and $\Lambda$ is the diagonal matrix of eigenvalues $\lambda_{1},\dots,\lambda_{n}$ of $L$. Any change to the eigenspectrum of $L$, clearly results in a change to $L_{rw}$, and therefore $P$. Since $\Psi_{0}=I-P$, any two graphs with distinct Laplacian eigenspectra will have distinct diffused dirac node embeddings. ∎
A.5 Detailed experimental settings
A complete overview of model hyperparameters and settings can be found in Table 4. Heuristically, the Graph-level MLP head hidden dimension is chosen to be the max # nodes multiplied by the hidden dimension size of the base GNN. We do NOT omit the trivial eigenvector when counting number of eigenvectors predicted.
Table 4: A comprehensive list of all model hyperparameters used during the eigenvector pre-training step. All hyperparameters highlighted in gray are specific to eigenvector-learning, while other listed configs reflect general GNN settings (and are set to match default values in each respective baseline work).
Method GIN (4.1) GIN (4.1) GIN pre-training (4.2) GPSE (4.3) Pre-training dataset ZINC-12k QM9 (134k) ZINC15 (2M), ChEMBL (456K) MolPCBA (324K) Base architecture GIN GIN GIN MPNN # params 33543 32883 2252210 22075899 # layers of per-node feature update 3 3 2 1 # layers of message passing 4 4 5 20 Hidden dim 60 60 300 512 Activation fn ReLU ReLU ReLU ReLU Dropout 0.1 0.1 0.2 0.2 Batch size 128 128 32 1024 Learning rate 0.001 0.001 0.001 0.005 Optimizer Adam Adam Adam AdamW Scheduler ReduceLROnPlateau ReduceLROnPlateau None CosineWithWarmup patience=5, factor=0.9 patience=20, factor=0.5 - - Pre-Training Epochs 200 100 100 120 Fine Tuning Epochs 500 150 100 - Laplacian norm type Unnormalized Unnormalized Unnormalized Symmetric # eigenvectors predicted 6 6 5 5 Initial features Diffused dirac + Wavelet pos. Diffused dirac + Wavelet pos. Molecule features Random MLP head type(s) Graph-level Graph-level Graph-level, per-node Graph-level, per-node Graph-level MLP max # nodes 40 40 50 50 MLP head # layers 5 5 1 2 MLP head hidden dim 2400 2400 N/A 1600, 32 MLP head activation fn ReLU ReLU N/A ReLU Loss function (and coefficient) 2*Eigenvector + 1*energy 2*Eigenvector + 1*energy 0.25 * Eigenvector + 0.05 * ortho 0.25 * Eigenvector Other features/notes Removed graphs with less than six nodes during pre-training Residual gating, virtual node
A.6 Modifying existing pre-training method
A.6.1 Learning rate tuning
We keep the majority of the settings from [3] the same. For downstream fine-tuning, we tune over 7 learning rates for fair comparison according to [4]. We run each method and learning rate over 3 seeds, and select the learning rate based on mean validation accuracy over all learning rates.
A.6.2 Downstream datasets
We briefly list and cite the five downstream datasets here for reference. The five datasets are the datasets chosen in [4], and are a subset of the eight primary downstream datasets evaluated in [3].
  • BACE: Qualitative binding results [19]
  • BBBP: Blood-brain barrier penetration [20]
  • Tox21: Toxicity data [21]
  • Toxcast: Toxicology measurements [22]
  • SIDER: Database of adverse drug reactions (ADR) [23]
A.7 Modifying existing graph structural encoder
A.7.1 Loss function comparison
We claim that the sum of MAE and cosine similarity loss used in [1] is ill-posed for learning eigenvectors.
To address sign ambiguity, the original GPSE model only learns the absolute value of each Laplacian eigenvector as its target. This improves performance over training without absolute value, but two shortcomings still arise. This (1) provides strictly less information than learning the true eigenvectors, and (2) does not capture basis ambiguity, making it ill-posed for eigenvectors with higher multiplicity (or eigenvectors with small eigengaps).
We demonstrate this claim experimentally in Figure 2. We train the default GPSE settings and architecture using (1) eigenvector loss (ours), and the default cosine similarity + MAE loss (GPSE). We train for 600 epochs on 1% of the MolPCBA dataset and plot eigenvector and energy losses per-epoch. We also plot the loss when outputting random orthogonal vectors for comparison. While training on cosine similarity + MAE achieves better-than-random results, sign and basis ambiguity issues severely limit the model’s ability to learn meaningful eigenvectors.
Figure 2: A comparison of energy loss and eigenvector loss, when the GPSE architecture is trained on (1) eigenvector loss (ours), (2) cosine similarity + MAE [1], and (3) no loss (outputting random vectors). All losses displayed are in natural log scale for visual purposes.